OpenStudy (anonymous):

The ratio of men to women in a certain factor is 3 to 4. There are 228 men. How many workers are there ? Check my answer please !

OpenStudy (anonymous):

304 workers was my guess

OpenStudy (a1234):

Proportion: 3/4 = 228/x You get 304 women. 304 + 228 = ?

OpenStudy (anonymous):

we're not looking for the amount of feamles. We're looking for the amount of workers from women to men combined 3 : 4 228 : ? 228/3 = 76 76 x 4 = 304

OpenStudy (a1234):

304 is the number of females. Remember, the ratio is men to women, not to workers. So add 304 and 228 to get 532.

OpenStudy (anonymous):

whoops so it'd be 532?

OpenStudy (a1234):

Yes, 532 workers.